# Quantitative periodic homogenization of parabolic equations with large drift and potential

### v1

@@ 1 Introduction
for some $\mu>0$. We will be commenting more on the initial datum $u^{0}_{\varepsilon}$ later.
@@
where $(\lambda,\psi)$ are an eigenpair associated with a certain exponential $Y$-periodic eigenvalue problem. Factorization similar to (1.2) are already available in [4, 3, 10]. The use of exponential $Y$-periodic eigenvalue problems in homogenization can be traced back to [6, page 383]. Allaire and Orive in [3, Lemma 2.1] (see also [7, Lemma 3], where Capdeboscq treats operators with no lower order terms) show the existence of a unique $\theta\in\mathbb{R}^{d}$ such that the eigenvalue problems (with $A^{\ast}$ denoting the adjoint of $A$):
@@
  • Both the problems (1.3) and (1.4) have the same first eigenvalue which is real, positive and simple.
  • The first eigenfunctions of (1.3) and (1.4) can be chosen such that
@@
The coefficient $\sigma$ is $Y$-periodic because
$\sigma(y)=\psi(y)\psi^{\ast}(y)=\psi(y)e^{-2\pi\theta\cdot y}e^{2\pi\theta\cdot y}\psi^{\ast}(y)$
Remark 1.2.
We have made a choice of taking the initial datum for $v_{\varepsilon}$ in (1.7) to be $\varepsilon$-independent. The factorization (1.2), then, forces the initial datum for $u_{\varepsilon}$ in (1.1) to be of the form
$u_{\varepsilon}(0,x)=u^{0}_{\varepsilon}(x)=u_{0}(x)\psi\left(\frac{x}{\varepsilon}\right)$
This is sometimes referred to as the well prepared initial datum. Similar well preparedness assumptions on the data have been made in the literature [2, 3]. This choice results in strong convergence of the factor $v_{\varepsilon}$ to its homogenized limit. If one were to keep the data $u_{\varepsilon}(0,x)$ independent of the parameter $\varepsilon$, then the factorization forces the initial datum for $v_{\varepsilon}$ to be
@@
All of this can be argued using the notion of two-scale convergence [1], for example. In this case, the convergence will be weak and hence cannot talk about rates.
Our main result is the following quantitative homogenization result for the factor $v_{\varepsilon}$ in (1.2) which solves the initial boundary value problem (1.7).
@@
Let $v_{\varepsilon}$ be the solution to (1.7) and let $v_{0}$ be the solution to the corresponding homogenized problem (1.12). Then, there exists a positive constant $C$, independent of $\varepsilon$, such that the following estimate holds:
$\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\varepsilon^{\frac{1}{4}}\left\{\left\|\partial_{t}v_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|v_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|u_{0}\right\|_{L^{2}(\Omega)}\right\}.$ (1.10)
In the above result, there is an inherent assumption that the terms on the right hand side of the bound are finite. This essentially translates to the data being sufficient smooth to guarantee adequate regularity for the solution $v_{0}$ of the homogenized equation. The proof of the above theorem is inspired by the ideas in the work of Jun Geng and Zhongwei Shen [11] that dealt with $\mathrm{L}^{2}$ convergence rates in the context of periodic homogenization of parabolic problems.
@@
Our estimate (1.10) is not sharp as in [11]. In there, the authors have obtained an $\mathcal{O}(\varepsilon)$ rate. Upon deriving a $\mathcal{O}(\sqrt{\varepsilon})$ rate for $\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$, the authors employ a duality argument to improve the rate to $\mathcal{O}(\varepsilon)$. We emphasise here that the same strategy can be followed in our case as well. Thus, improving the rate in Theorem 1.3 to $\mathcal{O}(\sqrt{\varepsilon})$.
Before we comment further on the proof of Theorem 1.3, we return to the equation (1.7). As $\beta$ (defined by (1.6)) is periodic, divergence-free and is of zero average, there exists a periodic skew-symmetric matrix-valued function $y\mapsto B(y)$ such that (see [7, Lemma 1, p. 809])
@@
$M(y)\xi\cdot\xi=\alpha(y)\xi\cdot\xi\geq b\left|\xi\right|^{2}\quad\forall\xi\in\mathbb{R}^{d}\,\mbox{ and for a.e. }y\in Y,$
where the ellipticity constant $b$ depends on the ellipticity constant $\mu$ of $A$ and the lower bound $a$ in (1.5). Using classical homogenization arguments (see [9, Theorem 11.4, p. 211]), we can show that
$\lim_{\varepsilon\to 0}\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}=0,$
@@
Factorizations comparable to (1.2) have appeared before [10, 4, 3]. We would like to emphasise here that [10] treats the setting where both the $\varepsilon$-problem and the corresponding homogenized problem are posed in the full space, i.e. the spatial domain is the whole of $\mathbb{R}^{d}$. The authors of [4] too are working in a perforated medium built out of the whole of $\mathbb{R}^{d}$. This is inherently due to the effective non-zero drift present in the model. In [3], the authors pose the problem in a bounded spatial domain $\Omega$. However, they make a change of unknown which takes the domain to the whole space $\mathbb{R}^{d}$ and incidentally the homogenized equation obtained in [3, Theorem 3.1] is posed in $\mathbb{R}^{d}\times[0,T)$. To the best of our knowledge, this present work is the first time where parabolic problems with large drift are treated and both the $\varepsilon$-problem and the homogenized problems are posed in $\Omega_{T}$ with $\Omega$ being a bounded domain in $\mathbb{R}^{d}$.
@@
The results of this paper also yields quantitative rates for the periodic homogenization of parabolic equations with nondivergence form elliptic part and with large lower order terms. We draw inspiration from the work of Wenjia Jing and Yiping Zhang [12] which derived quantitative rates for the periodic homogenization of nondivergence form elliptic equations with large drift terms. The key idea in [12] is to use the invariant measure associated with the underlying diffusion process. This is a well-known procedure (see [5, 6], also see [8, section 2.2]) to transform nondivergence form equations into divergence form. This present work deviates from the earlier works by not appealing to this invariant measure. We demonstrate that via exponential eigenvalue problems akin to (1.3)-(1.4), one can transform the parabolic equation with terms in nondivergence form into a parabolic equation with terms in divergence form. The initial-boundary value problem, for the unknown $\rho_{\varepsilon}:\overline{\Omega}\times[0,T)\to\mathbb{R}$, that we examine is the following:
@@
2 Quantitative rates in $\mathrm{L}^{2}(\Omega_{T})$
In this section, we prove a general quantitative homogenized result for the following parabolic problem:
@@ 2 Quantitative rates in $\mathrm{L}^{2}(\Omega_{T})$
In the above problem, the coefficient $\zeta\in\mathrm{L}^{\infty}_{\text{per}}(Y)$ and is assumed to satisfy
$\int_{Y}\zeta(y)\,{\rm d}y=1.$ (2.2)
The matrix-valued coefficient $\Theta\in\mathrm{L}^{\infty}_{\text{per}}(Y)$ is assumed to be uniformly elliptic, i.e.
@@
Our main result (Theorem 1.3) is a special case of the following quantitative result corresponding to the periodic homogenization of (2.1).
@@
Let $f_{\varepsilon}$ be the solution to (2.1) and let $f_{0}$ be the solution to the corresponding homogenized problem (2.4). Then, there exists a positive constant $C$, independent of $\varepsilon$, such that the following estimate holds:
@@
$\eta_{1}^{\varepsilon}=1\quad\mbox{ in }\quad\Omega\backslash\Omega_{4\sqrt{\varepsilon}},\qquad\eta_{1}^{\varepsilon}=0\quad\mbox{ in }\quad\Omega_{3\sqrt{\varepsilon}},\qquad\left|\nabla\eta_{1}^{\varepsilon}\right|\leq\frac{1}{\sqrt{\varepsilon}}\quad\mbox{ in }\quad\Omega.$
@@
$\eta_{2}^{\varepsilon}=1\quad\mbox{ in }\,\left(8\varepsilon,T-8\varepsilon\right),\qquad\eta_{2}^{\varepsilon}=0\quad\mbox{ in }\,(0,4\varepsilon]\cup(T-4\varepsilon,T),\qquad\left|\partial_{t}\eta_{2}^{\varepsilon}\right|\leq\frac{1}{\varepsilon}\quad\mbox{ in }\,(0,T).$
Next, we consider a parabolic smoothing operator borrowed from [11, Section 3, p.2102]:
@@
where $\theta\in\mathrm{C}^{\infty}_{c}(\mathcal{K})$ is a fixed nonnegative function that satisfies
@@
Here the set $\mathcal{K}\subset\mathbb{R}^{d+1}$ is defined below:
$\mathcal{K}:=\left\{(t,x)\in\mathbb{R}\times\mathbb{R}^{d}\mbox{ such that }\left|t\right|+\left|x\right|^{2}\leq 1\right\}.$
@@
where $f_{\varepsilon}$ and $f_{0}$ are solutions to (2.1) and (2.4) respectively. Note that $w_{\varepsilon}\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$ and that $w_{\varepsilon}(0,\cdot)\equiv 0$. We begin our analysis with the following result. In the remainder of this manuscript, we will sometimes be using the following shorthand notations:
@@
where we have used the fact that $f_{\varepsilon}$ and $f_{0}$ satisfy the evolution equations in (2.1) and (2.4), respectively. Owing to the anti-symmetric nature of the third-order tensor (i.e. $\phi_{kij}=-\phi_{ikj}$ for $1\leq i,j,k\leq d$), it follows that (see [13, Remark 3.1.3]) for a nice function $\psi$, we have
@@
$\leq C\varepsilon^{2}\Bigg\{\frac{1}{\varepsilon}\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\varepsilon}\left\|\frac{\partial f_{0}}{\partial t}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\sqrt{\varepsilon}}\left\|\frac{\partial f_{0}}{\partial t}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
Note that
$\frac{\partial^{2}}{\partial x_{i}\partial t}$ $\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]=S_{\varepsilon}\left(\frac{\partial^{2}}{\partial x_{i}\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial}{\partial x_{i}}\left(\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)\right)$
$=S_{\varepsilon}\left(\frac{\partial^{2}}{\partial x_{i}\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+\frac{\partial}{\partial x_{i}}\left(S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)\right)-S_{\varepsilon}\left(\frac{\partial}{\partial x_{i}}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)$
@@
$\hskip 184.9429pt+\left\|S_{\varepsilon}\Big(\nabla\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\partial_{t}\nabla f_{0}\Big)\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{2}\Bigg\{\frac{1}{\varepsilon^{\frac{3}{2}}}\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\varepsilon}\left\|\partial_{t}\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\sqrt{\varepsilon}}\left\|\partial_{t}\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
@@
$I_{2}\leq C\varepsilon^{\frac{1}{2}}\left\|f_{0}\right\|_{\mathrm{H}^{1}(\Omega_{T})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
@@
$\leq\left\|\nabla f_{0}-S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\|_{\mathrm{L}^{2}(\Omega_{T,4\sqrt{\varepsilon}})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T,4\sqrt{\varepsilon}})}$
$\hskip 170.71652pt+\left\|\nabla f_{0}-S_{\varepsilon}(\nabla f_{0})\right\|_{\mathrm{L}^{2}(\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}})}$

### v2

@@ 1 Introduction
for some $\mu>0$. We will be commenting more on the initial datum $u^{0}_{\varepsilon}$ later. In the remainder of this manuscript, the standard notation $\Omega_{T}$ is used to denote the parabolic cylinder.
@@
where $(\lambda,\psi)$ are an eigenpair associated with a certain exponential $Y$-periodic eigenvalue problem. The use of exponential $Y$-periodic eigenvalue problems in homogenization can be traced back to [7, page 383]. Allaire and Orive in [4, Lemma 2.1] (see also [8, Lemma 3], where Capdeboscq treats operators with no lower order terms) show the existence of a unique $\theta\in\mathbb{R}^{d}$ such that the eigenvalue problems (with $A^{\ast}$ denoting the adjoint of $A$):
@@
  • Both the problems (1.3) and (1.4) have the same first eigenvalue $(\lambda=\lambda^{\ast})$ which is real and simple.
  • The first eigenfunctions of (1.3) and (1.4) are bounded and can be chosen such that
@@
Note that the coefficient $\sigma$ is $Y$-periodic because
$\sigma(y)=\psi(y)\psi^{\ast}(y)=\psi(y)e^{-2\pi\theta\cdot y}e^{2\pi\theta\cdot y}\psi^{\ast}(y).$
@@
Our choice of taking the initial datum for $v_{\varepsilon}$ in (1.7) to be $\varepsilon$-independent forces, due to the factorization (1.2), the initial datum for $u_{\varepsilon}$ in (1.1) to be of the form:
$u_{\varepsilon}(0,x)=u^{0}_{\varepsilon}(x)=u_{0}(x)\psi\left(\frac{x}{\varepsilon}\right)$ (1.9)
This is sometimes referred to as the well prepared initial datum. Similar well preparedness assumptions on the data have been made in the literature [3, 4]. This choice results in strong convergence of the factor $v_{\varepsilon}$ to its homogenized limit. If one were to keep the data $u_{\varepsilon}(0,x)$ independent of the parameter $\varepsilon$, then the factorization (1.2) forces the initial datum for $v_{\varepsilon}$ to be
@@
All of this can be argued using the notion of two-scale convergence [1], for example. In this case, the convergence will be weak and hence one cannot talk about rates.
Our main result is the following quantitative homogenization result for the factor $v_{\varepsilon}$ which solves the initial boundary value problem (1.7):
@@
Let $v_{\varepsilon}$ be the solution to (1.7) and let $v_{0}$ be the solution to the corresponding homogenized problem (1.14). Then there exists a positive constant $C$, independent of $\varepsilon$, such that the following estimate holds:
$\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\sqrt{\varepsilon}\left\{\left\|\partial_{t}v_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|v_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|u_{0}\right\|_{L^{2}(\Omega)}\right\}.$ (1.11)
In the above result, there is an inherent assumption that the terms on the right hand side of the bound (1.11) are finite. This essentially translates to the data being sufficient smooth to guarantee adequate regularity for the solution $v_{0}$ of the homogenized equation. The proof of the above theorem is inspired by the ideas in the work of Jun Geng and Zhongwei Shen [12] that dealt with $\mathrm{L}^{2}$ convergence rates in the context of periodic homogenization of parabolic problems.
@@
Our estimate (1.11) is not sharp as in [12]. In there, the authors have obtained an $\mathrm{L}^{2}$ convergence rate of $\mathcal{O}(\varepsilon)$. Upon deriving a $\mathcal{O}(\sqrt{\varepsilon})$ rate for $\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$, the authors in [12] employ a duality argument to improve the rate to $\mathcal{O}(\varepsilon)$. We emphasise that the same strategy cannot be followed in our case as the bound (1.11) involves the $W^{1,d}(\Omega)$-norm. The matter of deriving optimal $\mathcal{O}(\varepsilon)$ rates for $\mathrm{L}^{2}$ convergence is left for future investigations. We comment that one may be able to employ the duality argument provided the following bound holds:
$\|\partial_{t}v_{0}\|_{\mathrm{L}^{2}(0,T;\mathrm{W}^{1,d}(\Omega))}\leq C\|f\|_{\mathrm{L}^{2}(\Omega_{T})}.$
Before we comment further on the proof of Theorem 1.2, we return to the equation (1.7). As $\beta$ (defined by (1.6)) is periodic, divergence-free and is of zero average, there exists a periodic skew-symmetric matrix-valued function $y\mapsto B(y)$ such that (see [8, Lemma 1, p. 809])
@@
$M(y)\xi\cdot\xi=\alpha(y)\xi\cdot\xi\geq\mathfrak{b}\left|\xi\right|^{2}\quad\forall\xi\in\mathbb{R}^{d}\,\mbox{ and for a.e. }y\in Y,$
where the ellipticity constant $\mathfrak{b}$ depends on the ellipticity constant $\mu$ of $A$ and the lower bound $a$ in (1.5). Using classical homogenization arguments (see [10, Theorem 11.4, p. 211]), we can show that
$\lim_{\varepsilon\to 0}\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}=0,$ (1.13)
@@
The above noted uniform ellipticity of $M$ helps us arrive at the following uniform (with respect to $\varepsilon$) estimate on the solution $v_{\varepsilon}$ to (1.12):
$\left\|v_{\varepsilon}\right\|_{\mathrm{L}^{\infty}([0,T];\mathrm{L}^{2}(\Omega))}\leq C\left\|u_{0}\right\|_{\mathrm{L}^{2}(\Omega)}.$ (1.16)
Appealing now to the factorization (1.2) we arrive at
$\left\|u_{\varepsilon}\right\|^{2}_{\mathrm{L}^{2}(\Omega_{T})}=\int_{0}^{T}\int_{\Omega}e^{-\frac{2\lambda t}{\varepsilon^{2}}}\left|\psi\left(\frac{x}{\varepsilon}\right)\right|^{2}\left|v_{\varepsilon}(t,x)\right|^{2}\,{\rm d}x\,{\rm d}t\leq C\int_{0}^{T}e^{-\frac{2\lambda t}{\varepsilon^{2}}}\,{\rm d}t,$ (1.17)
where we have used the boundedness of the eigenfunction $\psi$ in $\mathrm{L}^{\infty}$ and the uniform bound on $v_{\varepsilon}$ mentioned above. Next, subject to $\lambda$ being strictly positive, performing time integration in the bound (1.17) leads to
$\left\|u_{\varepsilon}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\leq C\varepsilon.$ (1.18)
As mentioned in Remark 1.1, in the absence of the well-prepared initial datum (1.9), the initial datum for $v_{\varepsilon}$ in (1.7) takes the form (1.10). Note that the uniform estimate (1.16) for $v_{\varepsilon}$ holds true even in this case. This implies that the above conditional (i.e. under the assumption of $\lambda>0$) decay rate (in $\varepsilon$) for the solution $u_{\varepsilon}$ to (1.1) remains true for general $\mathrm{L}^{2}$ initial data (i.e. no need for well-preparedness in the sense of (1.9)). More comments on the aforementioned conditional decay rate for $u_{\varepsilon}$ is relegated to our concluding remarks (see Section 4).
Remark 1.5.
@@
Factorizations comparable to (1.2) have appeared before [11, 5, 4]. We would like to emphasise here that [11] treats the setting where both the $\varepsilon$-problem and the corresponding homogenized problem are posed in the full space, i.e. the spatial domain is the whole of $\mathbb{R}^{d}$. The authors of [5] too are working in a perforated medium built out of the whole of $\mathbb{R}^{d}$. This is inherently due to the effective non-zero drift present in the model. In [4], the authors pose the problem in a bounded spatial domain $\Omega$. However, they make a change of unknown which takes the domain to the whole space $\mathbb{R}^{d}$ and incidentally the homogenized equation obtained in [4, Theorem 3.1] is posed in $\mathbb{R}^{d}\times[0,T)$. To the best of our knowledge, this present work is the first time where parabolic problems with large drift are treated with both the $\varepsilon$-problem and the homogenized problems are posed in $\Omega_{T}$ with $\Omega$ being a bounded domain in $\mathbb{R}^{d}$.
Remark 1.6.
Our analysis follows the approach of Geng and Shen [12] which relies on fine properties of a parabolic smoothing operator. If one builds parabolic boundary layer correctors, then the rates can be obtained without the involvement of these smoothing operators. This line of approach is inspired by the work of Allaire and Amar [2] which treated boundary layers in the context of elliptic problems. We will be addressing these points in a future publication.
@@
The results of this paper also yields quantitative rates for the periodic homogenization of parabolic equations with nondivergence form elliptic part and with large lower order terms. We draw inspiration from the work of Wenjia Jing and Yiping Zhang [13] which derived quantitative rates for the periodic homogenization of nondivergence form elliptic equations with large drift terms. The key idea in [13] is to use the invariant measure associated with the underlying diffusion process. This is a well-known procedure (see [6, 7], also see [9, section 2.2]) to transform nondivergence form equations into divergence form. This present work deviates from the earlier works by not appealing to this invariant measure. We demonstrate that via exponential eigenvalue problems akin to (1.3)-(1.4), one can transform the parabolic equation with terms in nondivergence form into a parabolic equation with terms in divergence form. The initial-boundary value problem, for the unknown $\rho_{\varepsilon}:\overline{\Omega}\times[0,T)\to\mathbb{R}$, that we examine is the following:
@@
Remark 1.7.
In the context of the quantitative periodic homogenization of nondivergence form elliptic equations, Jing and Zhang [13] left open the case involving potential terms of order $\mathcal{O}(\varepsilon^{-2})$. We point out that the present work does not give a definitive answer to that question. Nevertheless, we have treated the parabolic version of that case. The technique adapted in this work is tailored specifically for parabolic problems. We cite the work of Zhang [15] which treated the large potential case in divergence form (without the drift terms), but of $\mathcal{O}(\varepsilon^{-1})$.
We finish this rather long introduction by detailing the structure of the rest of the paper. Section 2 quantifies the convergence (1.13) for general second order parabolic problems with the positive coefficient next to the time derivative term being merely measurable and bounded. The main result of section 2 is Theorem 2.1 which gives a rate of $\mathcal{O}(\varepsilon^{\frac{1}{4}})$. Section 3 improves this rate to $\mathcal{O}(\varepsilon^{\frac{1}{2}})$ with an additional assumption that the coefficient of the time derivative is regular (see Theorem 3.1). The manuscript concludes with an appendix which gathers some useful results on the parabolic smoothing operator.
Acknowledgements
The author would like to thank his PhD advisor Harsha Hutridurga for introducing him to the exciting theory of homogenization and for his useful remarks during the preparation of this manuscript. The author has also greatly benefited from the discussions with Chandan Biswas and Debanjana Mitra on a preliminary version of this manuscript. The author acknowledges the support of NBHM-DAE (Government of India).
2 The case of merely bounded coefficients
In this and the next section, we will study the following general parabolic problem:
@@ 2 The case of merely bounded coefficients
We assume that the coefficient $\zeta\in\mathrm{L}^{\infty}_{\text{per}}(Y)$ satisfies
$\zeta(y)\geq c>0\,\mbox{ for a.e. }y\in Y\quad\mbox{ and }\quad\int_{Y}\zeta(y)\,{\rm d}y=1,$ (2.2)
where $c$ is a constant. In the next section, an additional regularity assumption is made on $\zeta$. The matrix-valued coefficient $\Theta\in\mathrm{L}^{\infty}_{\text{per}}(Y)$ is assumed to be uniformly elliptic, i.e.
@@
The following is a quantitative result corresponding to the periodic homogenization of (2.1).
@@
Suppose $\zeta\in\mathrm{L}_{per}^{\infty}(Y)$ satisfies (2.2). Let $f_{\varepsilon}$ be the solution to (2.1) and let $f_{0}$ be the solution to the corresponding homogenized problem (2.4). Then there exists a positive constant $C$, independent of $\varepsilon$, such that the following estimate holds:
@@
$\eta_{1}^{\varepsilon}=1\quad\mbox{ in }\quad\Omega\backslash\Omega_{4\sqrt{\varepsilon}},\qquad\eta_{1}^{\varepsilon}=0\quad\mbox{ in }\quad\Omega_{3\sqrt{\varepsilon}},\qquad\left|\nabla\eta_{1}^{\varepsilon}\right|\leq\frac{C}{\sqrt{\varepsilon}}\quad\mbox{ in }\quad\Omega.$
@@
$\eta_{2}^{\varepsilon}=1\quad\mbox{ in }\,\left(8\varepsilon,T-8\varepsilon\right),\qquad\eta_{2}^{\varepsilon}=0\quad\mbox{ in }\,(0,4\varepsilon]\cup(T-4\varepsilon,T),\qquad\left|\partial_{t}\eta_{2}^{\varepsilon}\right|\leq\frac{C}{\varepsilon}\quad\mbox{ in }\,(0,T).$
Next, we consider a parabolic smoothing operator borrowed from [12, Section 3, p.2102]:
@@
where $\theta\in\mathrm{C}^{\infty}_{c}(\mathcal{N})$ is a fixed nonnegative function that satisfies
@@
Here the set $\mathcal{N}\subset\mathbb{R}^{d+1}$ is defined below:
$\mathcal{N}:=\left\{(t,x)\in\mathbb{R}\times\mathbb{R}^{d}\mbox{ such that }\left|t\right|+\left|x\right|^{2}\leq 1\right\}.$
@@
where $f_{\varepsilon}$ and $f_{0}$ are solutions to (2.1) and (2.4) respectively. Note that $w_{\varepsilon}\in\mathrm{L}^{2}(0,T;\mathrm{H}^{1}_{0}(\Omega))$ and that $w_{\varepsilon}(0,\cdot)\equiv 0$. The study of a function defined similar to (2.7) is carried out in the quantitative analysis of periodic elliptic problems [14, Chapter 3, page 33]. We note, however, that the analysis carried out by Geng and Shen in [12, equation (1.14), page 2095] considers a function which also involves a term of order $\mathcal{O}(\varepsilon^{2})$. We will be employing such a construction in the next section. In the remainder of this manuscript, we will sometimes be using the following shorthand notations:
@@
where we have used the fact that $f_{\varepsilon}$ and $f_{0}$ satisfy the evolution equations in (2.1) and (2.4), respectively. Owing to the anti-symmetric nature of the third-order tensor (i.e. $\phi_{kij}=-\phi_{ikj}$ for $1\leq i,j,k\leq d$), it follows that (see [14, Remark 3.1.3]) for a nice function $\psi$, we have
@@
$\leq C\varepsilon^{2}\Bigg\{\frac{1}{\varepsilon}\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\varepsilon}\left\|\frac{\partial f_{0}}{\partial t}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\sqrt{\varepsilon}}\left\|\frac{\partial f_{0}}{\partial t}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})},$
where we have used the Lemma 4.1 from the Appendix. Note that
$\frac{\partial^{2}}{\partial x_{i}\partial t}$ $\left[S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial f_{0}}{\partial x_{j}}\right)\right]=S_{\varepsilon}\left(\frac{\partial^{2}}{\partial x_{i}\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+S_{\varepsilon}\left(\frac{\partial}{\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial^{2}f_{0}}{\partial x_{i}\partial x_{j}}\right)$
$\qquad\qquad\qquad\qquad+S_{\varepsilon}\left(\frac{\partial}{\partial x_{i}}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)+S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial}{\partial x_{i}}\left(\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)\right)$
$=S_{\varepsilon}\left(\frac{\partial^{2}}{\partial x_{i}\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial f_{0}}{\partial x_{j}}\right)+\frac{\partial}{\partial x_{i}}\left(S_{\varepsilon}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\frac{\partial^{2}f_{0}}{\partial t\partial x_{j}}\right)\right)+S_{\varepsilon}\left(\frac{\partial}{\partial t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\frac{\partial^{2}f_{0}}{\partial x_{i}\partial x_{j}}\right).$
@@
$\hskip 184.9429pt+\left\|S_{\varepsilon}\Big(\partial_{t}\left(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\right)\nabla^{2}f_{0}\Big)\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
$\leq C\varepsilon^{2}\Bigg\{\frac{1}{\varepsilon^{\frac{3}{2}}}\left\|\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\varepsilon}\left\|\partial_{t}\nabla f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}+\frac{1}{\varepsilon}\left\|\nabla^{2}f_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\Bigg\}\left\|\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}.$
@@
$I_{2}\leq C\sqrt{\varepsilon}\big\{\left\|\partial_{t}f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{W}^{1,d}(\Omega))}+\left\|f_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}\big\}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T})}$
@@
$\leq C\left\|\nabla f_{0}-S_{\varepsilon}(\eta_{1}^{\varepsilon}\eta_{2}^{\varepsilon}\nabla f_{0})\right\|_{\mathrm{L}^{2}(\Omega_{T,4\sqrt{\varepsilon}})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T,4\sqrt{\varepsilon}})}$
$\hskip 170.71652pt+C\left\|\nabla f_{0}-S_{\varepsilon}(\nabla f_{0})\right\|_{\mathrm{L}^{2}(\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}})}\left\|\nabla\psi\right\|_{\mathrm{L}^{2}(\Omega_{T}\setminus\Omega_{T,4\sqrt{\varepsilon}})}$